# RedChronos: A Large Language Model-Based Log Analysis System for Insider Threat Detection in Enterprises

### v1

@@ Abstract
Internal threat detection aims to address security threats within organizations or enterprises by identifying potential or already occurring malicious threats within vast amounts of logs. Although organizations or enterprises have dedicated personnel responsible for reviewing these logs, it is impossible to manually examine all logs entirely. In response to the vast number of logs, we propose a system called RedChronos, which is a Large Language Model-Based Log Analysis System. This system incorporates innovative improvements over previous research by employing Query-Aware Weighted Voting and a Semantic Expansion-based Genetic Algorithm with LLM-driven Mutations. On the public datasets CERT 4.2 and 5.2, RedChronos outperforms or matches existing approaches in terms of accuracy, precision, and detection rate. Moreover, RedChronos reduces the need for manual intervention in security log reviews by 90% in the Xiaohongshu SOC. Therefore, our RedChronos system demonstrates exceptional performance in handling Internal Threat Detection (IDT) tasks, providing innovative solutions for these challenges. We believe that future research can continue to enhance the system’s performance in IDT tasks while also reducing the response time to internal risk events.
@@ 4.2 Dataset
The internal SOC dataset from Xiaohongshu primarily consists of logs from various enterprise systems and security software. Logs from various components include email systems, gateway logs, and so on. The security software logs mainly originate from Xiaohongshu’s self-developed security software, Microsoft Defender for Endpoint (MDE), and Eagle Cloud Technology security logs. These logs contain detailed endpoint security information, including events such as employee device logins, device usage, potential risks from outdated operating systems, the presence of suspicious software, and indications of scanning or attacks on the machines.
@@ 5.1 Overall Performance
RedChronos 0.949 0.985 0.041 0.971
As shown in Table 1 and Table  2, RedChronos outperforms the state-of-the-art approach AuditLLM on public datasets in terms of Prec, DR, FPR, and Acc. Therefore, our research is currently considered state-of-the-art.
@@ 5.2 Ablation Study of Different Models Using the Same Prompt
DeepSeek R1 0.667 0.989 0.160 0.879
DeepSeek R1-Distill-Qwen-1.5B 0.464 0.729 0.270 0.729
DeepSeek R1-Distill-Qwen-7B 0.414 0.937 0.426 0.662
DeepSeek R1-Distill-Qwen-32B 0.662 0.991 0.163 0.876
Llama-3.3-70B-Instruct 0.332 0.991 0.644 0.512
DeepSeek-V3 0.933 0.987 0.022 0.979
As shown in Table 3, the performance of different LLM models is significantly impacted by using the same prompt for the final results. Therefore, it is essential to identify optimized prompts relevant to each specific task to maximize the model’s efficiency on that task.
@@ 6 Conclusion And Future Work
7 Appendix
We currently have several ongoing experiments, and we will continue to update the data and figures in the paper. The data presented in this paper represent a portion of the results from our research efforts.

### v2

@@ Abstract
Internal threat detection (IDT) aims to address security threats within organizations or enterprises by identifying potential or already occurring malicious threats within vast amounts of logs. Although organizations or enterprises have dedicated personnel responsible for reviewing these logs, it is impossible to manually examine all logs entirely. In response to the vast number of logs, we propose a system called RedChronos, which is a Large Language Model-Based Log Analysis System. This system incorporates innovative improvements over previous research by employing Query-Aware Weighted Voting and a Semantic Expansion-based Genetic Algorithm with LLM-driven Mutations. On the public datasets CERT 4.2 and 5.2, RedChronos outperforms or matches existing approaches in terms of accuracy, precision, and detection rate. Moreover, RedChronos reduces the need for manual intervention in security log reviews by approximately 90% in the Xiaohongshu Security Operation Center. Therefore, our RedChronos system demonstrates exceptional performance in handling IDT tasks, providing innovative solutions for these challenges. We believe that future research can continue to enhance the system’s performance in IDT tasks while also reducing the response time to internal risk events.
@@ 3.2.2 Semantic Expansion-based Genetic Algorithm
Since the algorithm requires us to iteratively refine the prompts and utilize a LLM to verify the effectiveness of our prompts, we also need a dataset to identify prompts with better performance. We selected 2000 abnormal and 2000 normal logs from the CERT v4.2 dataset in Sec 4.2.
We begin by mutating the Seed Prompt as demonstrated in Table 5. Subsequently, we leverage the LLM to evolve the Seed Prompt based on its performance outcomes, particularly analyzing both correct and incorrect response patterns. This iterative refinement process yields the optimized variations documented in Tables  6 and  7.
The first-generation prompts in Table 6 demonstrate suboptimal performance metrics, which consequently leads to their exclusion from subsequent evolutionary iterations through our selection mechanism. In contrast, the optimized prompt variants presented in Table 7 exhibit superior performance characteristics. Through five generations of systematic evolution with elitism preservation in Table 8, the final evolved prompt achieves satisfactory performance metrics.
@@ 4.1 Research Questions and Experimental Configuration
  3. 3.
    RQ3: Can RedChronos serve as an internal log analysis system for an enterprise or organization? How much daily workload can this system reduce?
@@ 4.2 Dataset
These logs contain detailed endpoint security information, including events such as employee device logins, device usage, potential risks from outdated operating systems, the presence of suspicious software, and indications of scanning or attacks on the machines.
@@
The internal SOC dataset from Xiaohongshu primarily consists of logs from various enterprise systems and security software. Logs from various components include email systems, gateway logs, and so on. The security software logs mainly originate from Xiaohongshu’s self-developed security software, Microsoft Defender for Endpoint (MDE), and Eagle Cloud Technology security logs. We use 10,000 MDE Alert logs from September 4, 2024, to November 4, 2024, as the dataset for this paper.
@@ 5.1 Overall Performance
RedChronos 0.975 0.993 0.019 0.988
As shown in Table 1 and Table  2, RedChronos outperforms the state-of-the-art approach AuditLLM on public datasets in terms of Prec, DR, FPR, and Acc. On the public datasets CERT 4.2 and 5.2, RedChronos outperforms or matches existing approaches in terms of accuracy, precision, and detection rate.
Superior DR and Acc. RedChronos achieves the highest DR (0.987 on r4.2, 0.993 on r5.2) and Acc (0.979 on r4.2, 0.988 on r5.2) across both datasets. This demonstrates its exceptional ability to identify true threats while maintaining high overall correctness.
Balanced Prec and FPR. While Audit-LLM marginally surpasses RedChronos in Precision on r4.2 (0.943 vs. 0.933), RedChronos achieves significantly lower FPR (0.022 vs. 0.037 on r4.2; 0.019 vs. 0.039 on r5.2). This highlights its robustness in minimizing false alarms—a critical factor for enterprise SOC efficiency.
Practical Relevance for SOC Workflows The extremely low FPR (0.022) and high Acc (0.979) across both datasets indicate RedChronos’ potential to drastically reduce manual review workloads (as claimed in Section 5.3) while ensuring reliable threat prioritization.
@@ 5.2 Ablation Study of Different Models Using the Same Prompt
DeepSeek-R1 [8] 0.667 0.989 0.160 0.879
DeepSeek-R1-Distill-Qwen-1.5B 0.464 0.729 0.270 0.729
DeepSeek-R1-Distill-Qwen-7B 0.414 0.937 0.426 0.662
DeepSeek-R1-Distill-Qwen-32B 0.662 0.991 0.163 0.876
Llama-3.3-70B-Instruct [14] 0.332 0.991 0.644 0.512
DeepSeek-V3 [9] 0.933 0.987 0.022 0.979
Table 4: On the CERT r5.2 dataset, the performance of various LLM models was evaluated using the same prompt.
Model CERT r5.2
Prec $\uparrow$ DR $\uparrow$ FPR $\downarrow$ Acc $\uparrow$
DeepSeek-R1 0.847 0.804 0.112 0.850
DeepSeek-R1-Distill-Qwen-1.5B 0.852 0.673 0.091 0.805
DeepSeek-R1-Distill-Qwen-7B 0.815 0.997 0.175 0.900
DeepSeek-R1-Distill-Qwen-32B 0.975 0.993 0.019 0.988
Llama-3.3-70B 0.552 0.999 0.631 0.645
DeepSeek-V3 0.949 0.985 0.041 0.971
As shown in Table 3 and  4, the performance of different LLM models is significantly impacted by using the same prompt for the final results. Tables 3 and  4 present a comparative evaluation of RedChronos against seven state-of-the-art methods on the CERT r4.2 and r5.2 datasets, respectively.
Model-Specific Performance Variability. Tables 3 and  4 reveal significant performance disparities across LLMs when using identical prompts. For instance, on CERT r4.2 (Table  3), DeepSeek-V3 achieves near-perfect detection rate (DR: 0.987) and accuracy (Acc: 0.979) with minimal false positives (FPR: 0.022), while Llama-3.3-70B-Instruct exhibits catastrophic FPR (0.644) despite high DR (0.991). This highlights the critical impact of model architecture and training strategies on task-specific performance. Smaller distilled models (e.g., Qwen-1.5B/7B) underperform in precision (0.464) and FPR (0.270), suggesting distillation sacrifices discriminative capability for efficiency.
Dataset-Specific Adaptation. Performance trends differ across CERT r4.2 and r5.2. For example, DeepSeek-R1-Distill-Qwen-32B shows improved precision on r5.2 (0.975 vs. 0.662 on r4.2), indicating better adaptation to r5.2’s threat patterns. Conversely, DeepSeek-R1 suffers a precision drop (0.667 to 0.847) but achieves higher DR (0.989 to 0.804) when transitioning from r4.2 to r5.2, implying a trade-off between specificity and sensitivity across datasets.
Hallucination vs. Robustness Trade-offs Models like Llama-3.3-70B achieve near-perfect DR (0.999 on r5.2) but at the cost of extreme FPR (0.631), reflecting hallucination tendencies. In contrast, DeepSeek-V3 maintains balanced performance (DR: 0.987–0.993, FPR: 0.019–0.022), demonstrating robustness against false alarms. This underscores the necessity of Query-Aware Weighted Voting to mitigate unreliable model outputs.
Scaling Benefits with Caveats. While larger models (e.g., Qwen-32B vs. Qwen-7B) generally improve precision and FPR, scaling alone is insufficient. For example, Qwen-32B outperforms its 7B counterpart in FPR (0.163 vs. 0.426 on r4.2), yet still lags behind DeepSeek-V3, emphasizing the need for specialized optimization (e.g., semantic-aware genetic algorithms) beyond mere parameter scaling.
Tables 3 and  4 underscore that LLM performance in IDT tasks is highly model- and dataset-dependent. RedChronos’s innovations in prompt optimization and ensemble voting address these variabilities, enabling reliable, low-intervention threat detection. Future work should explore dynamic model weighting based on real-time task feedback to further enhance adaptability.
5.3 Research on the Application of RedChronos in SOC
To address RQ 3, we have deployed RedChronos in Xiaohongshu’s SOC to assist security engineers and experts in log analysis. Based on historical log data in the SOC, RedChronos can reduce the number of log sample analyses performed by security engineers and experts by approximately 90%.